OpenStudy (anonymous):

x: 3(x + 1) = −2(x − 1) + 6 (earn a metal) 1 4 5 25

OpenStudy (anonymous):

x: 3(x + 1) = −2(x − 1) + 6 3x + 3 = -2x + 8 5x = 5 x = 1

OpenStudy (anonymous):

we first distribute the expressions on both sides of the equation. We then simplify each side of the equation, isolate and then solve for x